Aini and Bala had 2500 at first. After Aini spent 55, Bala had 4 times as much money as her. How much money did Aini have at first?
Answer by ikleyn(52864)

.
The problem says:

A + B = 2500,    (eq1)   and
4*(A-55) = B.    (eq2)


Substitute B = 4*(A-55)  from (eq2) into (eq1) to get 

A + 4*(A-55) = 2500.


Simplify and solve for A.   Then evaluate B.